If m is directly proportional to x and m = 2 when x = 12, then what is the value of x when m = 2^-1?
Answer by stanbon(75887)

If m is directly proportional to x and m = 2 when x = 12, then what is the value of x when m = 2^-1?
---------------
m = kx
---
Find "k"
2 = k*12
k = 1/6
--------------
So, m = (1/6)x
---
If m = 2^(-1) then
2^(-1) = (1/6)x
(1/2) = (1/6)x
Multiply both sides by 6 to solve for "x":
x = 3
=============================
Cheers,
Stan H.